# Enhancing Exploration with Diffusion Policies in Hybrid Off-Policy RL: Application to Non-Prehensile Manipulation

### v1

@@ Abstract
Learning diverse policies for non-prehensile manipulation is essential for improving skill transfer and generalization to out-of-distribution scenarios. In this work, we enhance exploration through a two-fold approach within a hybrid framework that tackles both discrete and continuous action spaces. First, we model the continuous motion parameter policy as a diffusion model, and second, we incorporate this into a maximum entropy reinforcement learning framework that unifies both the discrete and continuous components. The discrete action space, such as contact point selection, is optimized through Q-value function maximization, while the continuous part is guided by a diffusion-based policy. This hybrid approach leads to a principled objective, where the maximum entropy term is derived as a lower bound using structured variational inference. We propose the Hybrid Diffusion Policy algorithm (HyDo) and evaluate its performance on both simulation and zero-shot sim2real tasks. Our results show that HyDo encourages more diverse behavior policies, leading to significantly improved success rates across tasks - for example, increasing from 53% to 72% on a real-world 6D pose alignment task. Project page: https://leh2rng.github.io/hydo
@@ I INTRODUCTION
Figure 1: Overview of HyDO: The network takes point clouds, goal flow, and segmentation (indicating object and background points) as input. These are passed through the actor and critic networks. The actor is enhanced exploration on the continuous motion parameter with the entropy regularizer applied during the diffusion process and outputs the motion parameter. The state-action pair is then evaluated by the critic, which also integrates entropy regularization for exploration on the discrete contact location. The action with the highest Q-value is selected and executed by the robot.
@@ III-A MDP and Off-Policy Actor-Critic Methods
where $y_{t}$ is the target
$y_{t}=r_{t}+\gamma\mathbb{E}_{a_{t+1}\sim\pi_{\theta}(s_{t+1})}\left[Q_{\phi^{\prime}}(s_{t+1},a_{t+1})-\alpha\log\pi(a_{t+1}|s_{t+1})\right].$
@@ III-B Diffusion and Consistency Models
Consistency models [37] extend diffusion models by adopting the form of a probability flow ordinary differential equation (ODE). The reverse process along the ODE path $\{\hat{x}_{\tau}\}_{\tau\in[\epsilon,T]}$ generates data starting from $\hat{x}_{T}\sim\mathcal{N}(0,T^{2}I)$, where $\epsilon$ is a small value close to zero. The consistency model retains the effectiveness of a diffusion model but accelerates sampling by reducing the number of time steps.
@@ III-C Diffusion and Consistency Models as RL Policy
$\pi_{\theta}(\bm{a}|\bm{s}):=\pi_{\theta}(\bm{a}^{0:K}|\bm{s})={N}(\bm{a}^{K};\bm{0},\bm{I})\prod_{k=1}^{K}p_{\theta}(\bm{a}^{k-1}|\bm{a}^{k},\bm{s})$ (2)
@@ IV-A Problem Formulation
We formulate this problem in a principled manner as an online off-policy maximum entropy reinforcement learning task. This framework is chosen to encourage exploration and diversity in the learned behaviors. To represent the policies, we leverage Diffusion Probabilistic Models (DDPMs) [12] and Consistency Models (CMs) [37], which are well-suited for capturing diverse behaviors. Specifically, we first introduce a principled formulation to integrate diffusion policies into the Soft Actor-Critic (SAC) algorithm, enabling the continuous policies to capture multi-modalities. We then extend this approach to support both discrete and continuous actions within a hybrid RL framework, leading to our Hybrid Diffusion Policy (HyDo). This principled formulation enhances exploration across both action spaces, ensuring the development of diverse and robust manipulation policies.
@@ IV-B Soft Actor-Critic with Diffusion Policy
$\quad\left[\sum_{t=1}^{T}r(\bm{s}_{t},\bm{a}^{0}_{t})-\alpha\sum_{k=0}^{K}\log q(\bm{a}^{k-1}_{t}|\bm{a}^{k}_{t},k,\bm{s}_{t})\right].$
@@
$=\mathbb{E}_{\tau\sim q(\tau)}\left[\sum_{t=1}^{T}\left(r(\bm{s}_{t},\bm{a}_{t})-\alpha\sum_{k=1}^{K}\log q(\bm{a}_{t}^{k-1}|\bm{a}_{t}^{k},\bm{s}_{t})\right)\right].$
@@
$+\sum_{k=1}^{K}\nabla_{\bm{a}^{k-1}_{t}}\log p_{\theta}(\bm{a}^{k-1}_{t}|\bm{a}^{k}_{t},k,\bm{s}_{t})\nabla_{\theta}f_{\theta}(\epsilon^{k}_{t};\bm{a}^{k}_{t},k,\bm{s}_{t}),$
@@ IV-C Hybrid Diffusion Policy
In addition, the diffusion policy can be replaced by a consistency model $\pi^{m}(\bm{a}^{m}|\bm{s})=\texttt{Consistency\_Model}(\bm{s};f_{\theta})$ without changing the underlying optimization procedure (as shown in Algo. 1). We name this variant, Hybrid Diffusion Policy with Consistency Models (HyDo $+$ CM). Both variants are summarized with pseudo-code in Alg. 2.
Algorithm 1 Consistency Model Action Sampling
1:Given: state $\bm{s}$, $f_{\theta}$, a sub-sequence of time points $\{\tau_{n}\}_{n\in[N]}$, diffusion steps $K$.
2:Initialize mean $\bm{a}_{N}=\bm{0}$, variance $\Sigma_{N}=K^{2}\bm{I}$
3:for $n=N$ to 1 do
4:     Sample $\hat{\bm{a}}$ from ${N}(\bm{a}_{n},\Sigma_{n})$
5:     Compute $c_{\rm{skip}}=0.25/\left((\tau_{n}-\epsilon)^{2}+0.25\right)$ and $c_{\rm{out}}=0.5(\tau_{n}-\epsilon)/\sqrt{(\tau_{n}^{2}+0.25)}$
6:     Compute $\bm{a}_{n},\Sigma_{n}=f_{\theta}(\bm{s},\hat{\bm{a}},\tau_{n})$
7:     Compute output $\bm{a}_{n}=c_{\rm{skip}}\cdot\hat{\bm{a}}+c_{\rm{out}}\cdot\bm{a}_{n}$
8:end for
9:return $\bm{a}_{1}$
Algorithm 2 HyDo: Hybrid Diffusion Policy
3:Initialize replay buffer: ${D}=\emptyset$
4:while not converge do
5:     Forward the encoder to compute features $f=f(\bm{s}_{t})$
6:     Sample action map $\bm{a}^{m}_{t}$ from diffusion policy as in Eq. 3, e.g. DDPM or CM sampling presented in Algo. 1
7:     Compute Q-value map $Q=Q_{\phi}(f,\bm{a}^{m}_{t})$
8:     Select contact point $x_{i}$ using location policy Eq. 4
9:     Select corresponding action’s motion parameter $\bm{a}^{m}_{t,i}$
10:     Execute action $(x_{i},\bm{a}^{m,0}_{t,i})$, observe $r_{t}$ and $\bm{s}_{t+1}$
11:     Add sample $\{\bm{s}_{t},(x_{i},\bm{a}^{m,0}_{t,i}),r_{t},\bm{s}_{t+1}\}$ to replay buffer ${D}$
13:     Critic update as with target $y_{t}$ defined in IV-C.
14:     Actor update with loss $J_{\pi}(\theta)$ as defined in IV-C
15:     Adjust temperature $\alpha$
16:     Update the target networks like SAC.
17:end while
18:return final policy $\pi_{\theta}$.
@@ V Experiments
In this section, we evaluate our proposed algorithms, HyDo and HyDo $+$ CM, with the main baseline HACMan and its variants, HACMan $+$ Diff, and HACMan $+$ CM. In addition, we also investigate HyDo (w/o diffusion). All methods are evaluated on a set of simulated and real-world tasks. Our primary goal is to assess their performance in terms of success rate, behavior diversity, and generalization ability across different task settings. We use the same training settings as HACMan [7]. The input is a 4D point cloud obtained by concatenating 3D goal flow vectors with a 1D segmentation mask which indicates if the point belongs to the target object or the background. In simulation, ground-truth object masks are used as segmentation labels. In the real world robot experiments labels are obtained by background subtraction.
@@ V-A Experimental Setup
[…]
@@ V-B1 Simulation Results
Tab. I presents the evaluation results of the experiments conducted in simulation for 6D pose tasks on unseen category, unseen instance objects, as well as the objects in the training set. In this evaluation, we report the interquartile mean (IQM) [43] success rate over 10 seeds and using the best checkpoint for each method.
@@
Overall, these results indicate a substantial performance improvement with the introduction of the entropy term as a regularization. Specifically, HyDo and its variants consistently outperform the HACMan baselines on the Unseen Category and Unseen Instance object sets. The difference in performance is particularly highlighted in diffusion-based policies ($+$ Diff or $+$ CM), where HyDo and HyDo $+$ CM with the additional entropy regularization terms yields up to a 10% and 15% improvement compared to HACMan $+$ Diff and HACMan $+$ CM, respectively. In addition, either with or without the entropy terms, the diffusion policies achieve slightly better performance than consistency models across all evaluation sets.
@@ V-B2 Real Robot Results
We evaluate the trained policies on the ”All Objects + 6D Goals” simulation task on a real-world robot for which we use the same real-world setup, pose randomization, and success criteria as in HACMan (see Fig. 4).
(a) Lego
(b) Lotion
(c) Milk
(d) Soja
(e) Cube
Figure 4: Set of 5 objects used for real robot evaluations.
TABLE II: Real Robot Experiments for Planar (left) and 6D (right) Goals.
@@
Tab. II reports the evaluation results on both the planar and 6D tasks. We run each method with ten object trials and randomize the initial and target poses. To make it fair, we keep them similar across evaluations for all methods. The average success rates of the four methods are 53% for HACMan, 64% HyDo (w/o Diffusion), 68% for HyDo, and 72% for HyDo $+$ CM. These results demonstrate that all methods are also capable of effectively generalizing to unseen objects in real world. We do, however, observe a gap between non-diffusion and diffusion policy methods which is larger than in simulation. The gap between simulation and real world scenes is generally larger than the one for simulated training to simulated evaluation scenarios. Our hypothesis is that the diverse behaviors of the learned multi-modal diffusion policies allow for better generalization to such new environments. To further assess this diversity, we evaluate multi-modal action sequences under the same environment conditions, differing only by action sampling seeds. Fig. 2 and Fig. 3 show two distinct action sequences achieving the same goal pose in both simulation and real robot experiments, demonstrating how the order of flips and pushes can vary, achieving the same outcome through diverse execution actions.
@@ V-C Qualitative Policy Diversity Analysis
(a) Pushing task
(b) Entropy vs. Success Rates
Figure 5: Diverse behaviors of different diffusion-based policies on a simple pushing task (a). In (b), we show the Pareto plot between Entropy and Success Rates. Overall, methods without entropy regularization (HACMan) result in lower entropy compared to its counterparts. Note that we disable the entropy term on the location action to solely investigate the effect of the entropy on the continuous policies.
We follow the proposal to analyze the policy diversity as defined in [10]. They approximate behavior entropy of diffusion policies $\pi$ with the following function ${H}(\pi(\beta))=-\sum_{\beta\in{B}}\pi(\beta)\log_{|{B}|}\pi(\beta)$, where ${B}$ is defined as the set of task-specific behavior descriptors. We run simulations using the final policy of each method to estimate $\pi(\beta)$. We design a controlled task with $|{B}|=2$ where the robot can push the target object (on the left) along the upper or below path around the white static obstacle object to the target pose (on the right) as depicted in Fig. 5(a). If the entropy $H(\pi(\beta))=0$, it indicates that the policy has collapsed to a single solution or a single path (either upper or lower).
We qualitatively evaluate only methods with policies represented by diffusion and consistency models. More specifically, we train HyDo, HyDo $+$ CM, HACMan $+$ Diffusion, and HACMan $+$ CM until convergence with random initial poses. Then, we evaluate them for 200 trials with the same initial pose depicted in Fig. 5(a). In this experiment, to highlight the effect of the entropy on the continuous motion parameter, we disable the entropy term for the location action. As shown in Fig. 5(b), methods without entropy regularization such as HACMan, and HACMan $+$ CM result in less diverse behaviors (low entropy) and/or non-optimal actions (low success rate).
V-D Computational Efficiency Analysis of Consistency and Diffusion Models
To assess the computational speed of consistency and diffusion models with varying denoising steps, we run an experiment measuring inference times for $K\in\{2,5,10,20,50\}$, using a single-object ”Hammer” training setup. As shown in Tab. LABEL:tab:computation_time, the consistency model [37] achieves similar performance to the diffusion model with fewer denoising steps, demonstrating its efficiency advantage. Both models reach a performance plateau around $K=5$, with further increases in $K=50$ leading to a drop in performance.
TABLE III: Inference Time in milliseconds per sample for each method and diffusion step. (using IQM with 95% confidence intervals)

### v2

@@ Abstract
Learning diverse policies for non-prehensile manipulation is essential for improving skill transfer and generalization to out-of-distribution scenarios. In this work, we enhance exploration through a two-fold approach within a hybrid framework that tackles both discrete and continuous action spaces. First, we model the continuous motion parameter policy as a diffusion model, and second, we incorporate this into a maximum entropy reinforcement learning framework that unifies both the discrete and continuous components. The discrete action space, such as contact point selection, is optimized through Q-value function maximization, while the continuous part is guided by a diffusion-based policy. This hybrid approach leads to a principled objective, where the maximum entropy term is derived as a lower bound using structured variational inference. We propose the Hybrid Diffusion Policy algorithm (HyDo) and evaluate its performance on both simulation and zero-shot sim2real tasks. Our results show that HyDo encourages more diverse behavior policies, leading to significantly improved success rates across tasks - for example, increasing from 53% to 72% on a real-world 6D pose alignment task. Project page is available at https://leh2rng.github.io/hydo
Index Terms:
Reinforcement Learning; Machine Learning for Robot Control; Dexterous Manipulation
@@ I Introduction
Figure 1: Overview of HyDo: The network takes point clouds, goal flow, and segmentation (indicating object and background points) as input. These are passed through the actor and critic networks. The actor is enhanced exploration on the continuous motion parameter with the entropy regularizer applied during the diffusion process and outputs the motion parameter. The state-action pair is then evaluated by the critic, which also integrates entropy regularization for exploration on the discrete contact location. The action parameter and contact location with the highest Q-value is selected and executed by the robot.
@@ III-A MDP and Off-Policy Actor-Critic Methods
where $y_{t}$ is the target value, described as
$y_{t}=r_{t}+\gamma\,\mathbb{E}_{a_{t+1}\sim\pi_{\theta}(s_{t+1})}\big{[}$ $Q_{\phi^{\prime}}(s_{t+1},a_{t+1})$
$-\alpha\log\pi(a_{t+1}|s_{t+1})\big{]}.$
@@ III-B Diffusion and Consistency Models
However, diffusion models rely on progressive denoising over a large number of steps, which can lead to slow sampling speed. This can be a critical bottleneck for online RL where it heavily depends on sampling from environments. Consistency models [37] extend diffusion models by adopting the form of a probability flow ordinary differential equation (ODE). The reverse process along the ODE path $\{\hat{x}_{\tau}\}_{\tau\in[\epsilon,T]}$ generates data starting from $\hat{x}_{T}\sim\mathcal{N}(0,T^{2}I)$, where $\epsilon$ is a small value close to zero. Consistency models learn a parameterized function $f_{\theta}:(x_{\tau},\tau)\rightarrow x_{\epsilon}$ which maps from a noisy sample $x_{\tau}$ at step $\tau$ to its original sample $x_{\epsilon}$. Furthermore, as detailed in Section V-D, our experiments show that, despite a modest speedup, consistency models empirically achieve a higher success rate than pure diffusion methods.
@@ III-C Diffusion and Consistency Models as RL Policy
$\pi_{\theta}(\bm{a}|\bm{s}):=\pi_{\theta}(\bm{a}^{0:K}|\bm{s})=\mathcal{N}(\bm{a}^{K};\bm{0},\bm{I})\prod_{k=1}^{K}p_{\theta}(\bm{a}^{k-1}|\bm{a}^{k},\bm{s})$ (2)
@@ IV-A Problem Formulation
We formulate this problem in a principled manner as an online off-policy maximum entropy reinforcement learning task. This framework is chosen to encourage exploration and diversity in the learned behaviors. To represent the policies, we leverage Denoising Diffusion Probabilistic Models (DDPMs) [12] and Consistency Models (CMs) [37], which are well-suited for capturing diverse behaviors. Specifically, we first introduce a principled formulation to integrate diffusion policies into the Soft Actor-Critic (SAC) algorithm, enabling the continuous policies to capture multi-modalities. We then extend this approach to support both discrete and continuous actions within a hybrid RL framework, leading to our Hybrid Diffusion Policy (HyDo). This principled formulation enhances exploration across both action spaces, ensuring the development of diverse and robust manipulation policies.
@@ IV-B Soft Actor-Critic with Diffusion Policy
$\quad\left[\sum_{t=1}^{T}r(\bm{s}_{t},\bm{a}^{0}_{t})-\alpha\sum_{k=0}^{K}\log q(\bm{a}^{k-1}_{t}|\bm{a}^{k}_{t},k,\bm{s}_{t})\right],$
@@
$=\mathbb{E}_{\tau\sim q(\tau)}\left[\sum_{t=1}^{T}\left(r(\bm{s}_{t},\bm{a}_{t})-\alpha\sum_{k=1}^{K}\log q(\bm{a}_{t}^{k-1}|\bm{a}_{t}^{k},\bm{s}_{t})\right)\right],$
@@
$+\sum_{k=1}^{K}\nabla_{\bm{a}^{k-1}_{t}}\log p_{\theta}(\bm{a}^{k-1}_{t}|\bm{a}^{k}_{t},k,\bm{s}_{t})$
$\quad\cdot\nabla_{\theta}f_{\theta}(\epsilon^{k}_{t};\bm{a}^{k}_{t},k,\bm{s}_{t})$
@@ IV-C Hybrid Diffusion Policy
Algorithm 1 HyDo: Hybrid Diffusion Policy
3:Initialize replay buffer: ${D}=\emptyset$
4:while not converge do
5:     Forward the encoder to compute features $f=f(\bm{s}_{t})$
6:     Sample action map $\bm{a}^{m}_{t}$ from diffusion policy as in Eq. 2, e.g. DDPM or CM sampling.
7:     Compute Q-value map $Q=Q_{\phi}(f,\bm{a}^{m}_{t})$
8:     Select contact point $x_{i}$ using location policy Eq. 3
9:     Select corresponding action’s motion parameter $\bm{a}^{m}_{t,i}$
10:     Execute action $(x_{i},\bm{a}^{m,0}_{t,i})$, observe $r_{t}$ and $\bm{s}_{t+1}$
11:     Add $\{\bm{s}_{t},(x_{i},\bm{a}^{m,0}_{t,i}),r_{t},\bm{s}_{t+1}\}$ to replay buffer ${D}$
13:     Critic update as with target $y_{t}$ defined in IV-C.
14:     Actor update with loss $J_{\pi}(\theta)$ as defined in IV-C
15:     Adjust temperature $\alpha$
16:     Update the target networks like SAC.
17:end while
18:return final policy $\pi_{\theta}$.
@@
In addition, the diffusion policy can be replaced by a consistency model $\pi^{m}(\bm{a}^{m}|\bm{s})=\texttt{Consistency\_Model}(\bm{s};f_{\theta})$ without changing the underlying optimization procedure. We name this variant, Hybrid Diffusion Policy with Consistency Models (HyDo $+$ CM). Both variants are summarized with pseudo-code in Alg. 1.
@@ V Experiments
In this section, we evaluate HyDo and HyDo $+$ CM, and compare them with the main baseline HACMan and its variants, HACMan $+$ Diff, and HACMan $+$ CM. In addition, we also investigate HyDo (w/o diffusion). All methods are evaluated on a set of simulated and real-world tasks. Our primary goal is to assess their performance in terms of success rate, behavior diversity, and generalization ability across different task settings. We use the same training settings as HACMan [7]. The input is a 4D point cloud obtained by concatenating 3D goal flow vectors with a 1D segmentation mask which indicates if the point belongs to the target object or the background. Segmentation masks are obtained from ground-truth object masks (simulation) or by using background subtraction (real world robot experiments).
@@ V-A Experimental Setup
(a) Lego
(b) Lotion
(c) Milk
(d) Soja
(e) Cube
Figure 2: Set of five objects used for real robot evaluations.
@@ V-B1 Simulation Results
Tab. I presents the results of the experiments conducted in simulation for 6D pose tasks on unseen category, unseen instance objects, as well as the objects in the training set. In this evaluation, we report the interquartile mean (IQM) [43] success rate over 10 seeds and using the best checkpoint for each method.
@@
Overall, these results indicate a substantial performance improvement for methods which use an entropy term for regularization. Specifically, HyDo and its variants consistently outperform the HACMan baselines on the Unseen Category and Unseen Instance object sets. The difference in performance is particularly highlighted in diffusion-based policies ($+$ Diff or $+$ CM), where HyDo and HyDo $+$ CM with the additional entropy regularization terms yields up to a 10% and 15% improvement compared to HACMan $+$ Diff and HACMan $+$ CM, respectively. In addition, either with or without the entropy terms, the diffusion policies achieve slightly better performance than consistency models across all evaluation sets.
@@ V-B2 Real Robot Results
We follow the same setup, e.g., pose randomization, and success criteria as described in HACMan using a set of objects (see Fig. 2) to evaluate the “All Objects + 6D Goals” on the real-world robot setup, using the trained policies in simulation.
TABLE II: Success rates for real robot experiments on
Planar (left) and 6D (right) goals.
@@
Tab. II reports the evaluation results on both the planar and 6D tasks. We run each method with ten trials and randomize the initial and target object poses. We keep the initial and target poses similar across evaluations for all methods. The average success rates of the four methods are 53% for HACMan, 64% for HyDo (w/o Diffusion), 68% for HyDo, and 72% for HyDo $+$ CM. While these results demonstrate that all methods are capable of generalizing to unseen objects in real world scenarios, we do observe a gap between non-diffusion and diffusion policy methods. Compared to an entirely simulated evaluation, this gap is larger if the corresponding method is trained in simulation and applied to real world scenes. Our hypothesis is that the diverse behaviors of the learned multi-modal diffusion policies allow for better generalization to such new environments. To further assess this diversity, we evaluate multi-modal action sequences under the same environmental conditions, differing only by action sampling seeds. Fig. 3 shows two distinct action sequences achieving the same goal pose in real robot experiments. This demonstrates how the order of flips and pushes can vary through a diverse execution of actions while it still achieves the same outcome.
@@ V-C Qualitative Policy Diversity Analysis
Method Entropy
HACMan 0.000
HACMan + Diff 0.432
HACMan + CM 0.462
HyDo (w/o Diff) 0.000
HyDo 0.598
HyDo + CM 0.461
Figure 4: Left: A pushing task with $24$ behavioral modes. Right: Behavior entropy of different methodologies.
Figure 5: Left: A push and align task with $2$ behavioral modes. Right: Pareto plot between Entropy and Success Rates.
Due to the intractability of the diffusion policy’s entropy $H(\pi_{\theta})$, we follow (Jia et al. 2024) [10] to analyze policy diversity via behavior entropy, defined as
${H}(\pi(\beta))=-\sum_{\beta\in{B}}\pi(\beta)\log_{|{B}|}\pi(\beta),$
where ${B}$ is a set of task-specific behavior descriptors. We first evaluate our method on a task similar to [10], where the robot pushes a cube across a red line (see Fig. 4, left), with $|{B}|=24$ denoting the number of possible behaviors. We also introduce a new task where the robot must push and align the cube to a target pose by following either the upper or lower trajectory while avoiding white obstacles (see Fig. 5, left), resulting in $|{B}|=2$.
To enhance diversity during training, we randomly initialize the cube’s pose and train all methods until convergence. In the evaluation phase, we fix the initial pose (illustrated in Fig. 4 and Fig. 5), and perform 500 and 200 trials for pushing and the push-and-align tasks, respectively. In addition, the entropy term for the location actions is disabled to highlight the effect of the entropy on the continuous motion parameters.
Figure 6: Cluttered environment. The manipulated object is shown in color, the obscured object is the target, and the scene objects are depicted in dark wood.
In the pushing task, HACMan and HyDo (w/o Diff) fail to explore alternative strategies and converge to a single solution (Fig. 4, right), while diffusion-based methods solve the task using diverse trajectories. For the push-and-align task (Fig. 5, right), methods without entropy regularization (e.g., HACMan and HACMan $+$ CM) show reduced behavior diversity (low entropy) and/or lower success rates.
Next, we introduce the cluttered environment featuring varying obstacle objects (represented as a scene point cloud), as shown in Fig. 6. Table III presents the results, indicating that HyDo and HyDo $+$ CM outperform the baselines.
TABLE III: Performance under different cluttered scenes.
Method Numbers of Scene Objects
1 2 5
HACMan $0.527\pm 0.212$ $0.495\pm 0.155$ $0.421\pm 0.176$
HyDo (w/o Diff) $0.486\pm 0.121$ $0.457\pm 0.113$ $0.350\pm 0.110$
HyDo $0.601\pm 0.100$ $\mathbf{0.592\pm 0.091}$ $\mathbf{0.499\pm 0.080}$
HyDo + CM $\mathbf{0.615\pm 0.046}$ $0.585\pm 0.059$ $0.488\pm 0.042$
V-D Ablation Study of Consistency and Diffusion Models
To assess the computational speed of consistency and diffusion models with varying denoising steps, we run an experiment measuring inference times for $K\in\{5,10,20,50\}$, using a single-object ”Hammer” training setup. As shown in Tab. LABEL:tab:computation_time, the consistency model [37] achieves similar performance to the diffusion model with fewer denoising steps, demonstrating its efficiency advantage. Both models reach a performance plateau around $K=5$, with further increases in $K=50$ leading to a drop in performance.
TABLE IV: Inference Time in milliseconds per sample
for each method and diffusion step.